OpenStudy (anonymous):

Can somebody explain the mechanic of why the derivative of x^x = x^x(ln(x) + 1) ...thanks

OpenStudy (anonymous):

let's just do the right side x^x----> x ln(x) 1+ ln(x)

OpenStudy (anonymous):

you would want to use logarithmic differentiation on:\[y=x^x\]like we did in that last problem you asked.

OpenStudy (anonymous):

the derivative of x^x is really neat, it is like a hybrid of the exponential rule: c^x --> c^x(lnc) and the power rule: x^n --> nx^(n-1) because if you distribute the answer above: x^x(lnx + 1), you get x^x * lnx + x^x If you were to simply treat x^x like an exponential function, you would get the first term x^x * lnx and if you were to treat x^x like a power function, you get the second term because n is x: x^x --> x*x^(x-1) = x^x It may be misleading to point this out, but I find it reassuring.

OpenStudy (anonymous):

Of course the way you get it isn't by adding the two at all, it is by logarithmic differentiation, start with y = x^x and take the natural log of both sides: lny = lnx^x use rules of logs: lny = xlnx take derivative, using the chain rule on the left and the product rule on the right: y'/y = x*1/x + lnx simplify: y'/y = 1 + lnx multiply through by y, which is x^x: y' = x^x(1+lnx)

OpenStudy (anonymous):

the meaning of \[x^x\] is \[e^{x\ln(x)}\] now take the derivative using the chain rule. all the work is in the exponent, since the derivative of \[e^{\text{ something }}\] is \[e^ {\text{ something }}\times \frac{d}{dx}\text{ something}\]